OpenStudy (anonymous):

solve. 2log4-log3+2logx-4=0

OpenStudy (anonymous):

2log4-log3+2logx-4=0 Equivalent to \[\log(16/3)+\log(x ^{2})=4\] then \[\log(16x ^{2}/3)=4\] you will get \[x^{2}=10^{4}\times3\div16\] solve using calculator you will get 43.3 (1 decimal)